OpenStudy (anonymous):

The volume of a cube is 54 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@AliceAlexis

OpenStudy (anonymous):

@xostarbucks

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

a cube has all 3 dimensions equal what is the formula for the volume of the pyramid?

OpenStudy (anonymous):

V = 1 over 3(B)(h)

OpenStudy (triciaal):

V = 1/3*Bh where B is the base area

OpenStudy (anonymous):

im not sure

OpenStudy (triciaal):

volume of pyramid is 1/3 the volume of the cube because wow i can explain this one B = length * width and because the cube fits then the length and the width are equal the height is also the same as the cube since the cube fits the volume of pyramid is 1/3 *(l^2)*l = 1/3*l^3 volume of the cube is l^3 for a cube all the sides are equal l = length

OpenStudy (triciaal):

correction the pyramid fits inside the cube

OpenStudy (triciaal):

volume of pyramid = 1/3 volume of cube = 1/3*54 = 18

OpenStudy (triciaal):